OpenStudy (anonymous):

evaluate integral x^3 ln x dx

OpenStudy (allank):

Integration by parts would work nicely here.

OpenStudy (allank):

Stepwise:

OpenStudy (allank):

integral x^3 ln x dx 1) Differentiate ln x = 1/x ; Integrate x^3 = (x^4)/4 2) Rewrite integral expression integral x^3 ln x dx = (lnx)(x^4)/4 - integral ((x^4)/4)(1/x) =(x^4lnx)/4 - x^4/16 + constant